# WorldMem: Long-term Consistent World Simulation with Memory

### v1

@@ 4 Experiments
Datasets. We use MineDojo [6] to create diverse training and evaluation datasets in Minecraft, configuring diverse environments (e.g., plains, savannas, ice plains, and deserts), agent actions, and interactions. For real-world scenes, we utilize RealEstate10K [51] with camera pose annotations, designing trajectories that incorporate past scenes to evaluate long-term world consistency.
Metrics. For quantitative evaluation, we employ reconstruction metrics, where the method of obtaining ground truth (GT) varies by specific settings. We then assess the consistency and quality of the generated videos using PSNR, LPIPS [49], and reconstruction FID (rFID) [16], which collectively measure pixel-level fidelity, perceptual similarity, and overall realism.
@@ 4.1 Results on Generation Benchmark
Comparisons on Minecraft Benchmark. We compare our approach with a standard full-sequence training method [14, 40] and Diffusion Forcing (DF) [3]. The key differences are as follows: the full-sequence conditional diffusion transformer [28] maintains the same noise level during training and inference, DF introduces different noise levels for training and inference, and our method incorporates a memory mechanism. To assess both short-term and long-term world consistency, we conduct evaluations within and beyond the context window. We evaluate both settings on 300 test videos and use 4,800 frames to compute rFID. In the following experiments, the agent’s poses are generated by the game simulator as ground truth. However, in real-world scenarios, only the action input is available, and the pose is not directly observable. In such cases, the next-frame pose can be predicted based on the previous scenes, past states, and the upcoming action. We explore this design choice in the supplementary material.
@@
Figure 5: Within context window evaluation examples. It illustrates an example where the motion sequence first involves turning right and then returning to the original position, demonstrating methods’ ability to maintain self-contained consistency.
Table 1: Evaluation on Minecraft benchmark [6]
Within context window
Methods PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Full Sequence 20.35 0.0691 13.87
Diffusion Forcing [3] 26.56 0.0094 13.88
Ours 27.01 0.0072 13.73
Beyond context window
Methods PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Full Sequence / / /
Diffusion Forcing [3] 18.04 0.4376 51.28
Ours 25.32 0.1429 15.37
Figure 6: Beyond context window evaluation examples. It shows that Diffusion-Forcing suffers from inconsistency and quality degradation after generating a certain number of frames. In contrast, our method maintains high quality and faithfully reconstructs previously observed scenarios.
Comparisons on Real Scenarios. We further compare our approach with DFoT [32] on the RealEstate dataset [51]. Similar to DF [3], DFoT discards previously generated content that lies outside its context window and fails to maintain global world consistency.
To verify this, we currate 100 samples and test the 360-degree consistency of both DFoT and our method by rotating the scene a full 360 degrees and checking whether it returns to the same position. We measure the quality of reconstruction by comparing the first frame and the final frame (after a full rotation). As shown in Table 2, our method outperforms DFoT in all metrics, indicating that our integrated memory mechanism substantially improves consistency. Qualitative comparisons in Figure 7 further support these findings.
Figure 7: Examples on RealEstate [51]. DFoT [32] discards content beyond its context window, losing 360-degree consistency. In contrast, our method preserves details and accurately returns to the original location.
Table 2: Evaluation on RealEstate10K [51]
Methods PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
DFoT [32] 8.396 0.6676 156.74
Ours 20.19 0.1773 67.14
Qualitative results. Figure 4 showcases WorldMem’s capabilities. The top section demonstrates its ability to operate in a free action space across diverse environments. Given a 600-frame memory bank, our model generates 100 future frames while preserving the ground truth’s actions and poses, ensuring strong world consistency. The bottom section highlights dynamic environment interaction. By using timestamps as embeddings, the model remembers environmental changes and captures natural event evolution, such as plant growth over time.
@@ 4.2 Ablation
Embedding designs. The design of embeddings within the memory block is crucial for effective cross-frame relationship modeling. We evaluate three strategies (Table 3): (1) sparse pose embedding with absolute encoding, (2) dense pose embedding with absolute encoding, and (3) dense pose embedding with relative encoding. Results show that dense pose embeddings (Plücker embedding) significantly enhance all metrics, emphasizing the benefits of richer pose representations. Switching from absolute to relative encoding further improves performance, particularly in LPIPS and rFID, by facilitating relationship reasoning and information retrieval. As illustrated in Figure 8, absolute embeddings accumulate errors over time, while relative embeddings maintain stability even beyond 300 frames.
Table 3: Ablation on embedding designs
Pose type Embed. type PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Sparse Absolute 20.67 0.2887 39.23
Dense Absolute 23.63 0.1830 29.34
Dense Relative 25.32 0.1429 15.37
@@
Time condition. We ablate the effectiveness of timestamp embedding in Table 4. For this experiment, we curate 100 video samples featuring placing events and evaluate whether future generations align with event progression. As shown in the table, incorporating time embeddings significantly improves PSNR and LPIPS, indicating that adding temporal information helps the model faithfully reproduce event changes in world simulation. Since events like plant growth are inherently unpredictable, we do not conduct quantitative evaluations on such cases but instead provide qualitative illustrations in Figure 9.
@@
Time embedding PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
w/o 23.17 0.1989 23.89
w/ 25.12 0.1613 16.53
Memory retrieve strategy. We analyze memory retrieval strategies in Table 5. Random sampling from the memory bank leads to poor performance and severe quality degradation, as evidenced by a sharp drop in rFID and rapid divergence from the ground truth (Figure 8). The confidence-based filtering significantly enhances consistency and generation quality. Additionally, we refine retrieval by filtering out redundant memory units based on similarity, further improving all evaluation metrics and demonstrating the effectiveness of our approach.
Table 5: Ablation on memory retrieve strategy
Strategy PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Random 18.32 0.3224 47.35
+ Confidence Filter 23.12 0.1863 24.33
+ Similarity Filter 25.32 0.1429 15.37
@@ 5 Conclusion
In conclusion, WorldMem tackles the longstanding challenge of maintaining long-term consistency in world simulation by employing a memory bank of past frames and associated states. Its memory attention mechanism enables accurate reconstruction of previously observed scenes – even under large viewpoints or temporal gaps – and effectively models dynamic changes over time. Extensive experiments in both virtual and real settings confirm WorldMem’s capacity for robust, immersive world simulation. We hope our work will encourage further research on the design and applications of memory-based world simulators.

### v2

@@ 4 Experiments
Figure 4: Within context window evaluation. The motion sequence involves turning right and returning to the original position, showing self-contained consistency.
Figure 5: Beyond context window evaluation. Diffusion-Forcing suffers inconsistency over time, while ours maintains quality and recovers past scenes.
Table 1: Evaluation on Minecraft
Within context window
Methods PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Full Seq. 14.35 0.0691 13.87
DF 20.56 0.0094 13.88
Ours 21.01 0.0072 13.73
Beyond context window
Methods PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Full Seq. / / /
DF 18.04 0.4376 51.28
Ours 19.32 0.1429 15.37
Table 2: Ablation on embedding designs
Pose type Embed. type PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Sparse Absolute 14.67 0.2887 39.23
Dense Absolute 17.63 0.1830 29.34
Dense Relative 19.32 0.1429 15.37
Table 3: Ablation on memory retrieve strategy
Strategy PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Random 12.32 0.3224 47.35
+ Confidence Filter 17.12 0.1863 24.33
+ Similarity Filter 19.32 0.1429 15.37
Datasets. We use MineDojo (fan2022minedojo) to create diverse training and evaluation datasets in Minecraft, configuring diverse environments (e.g., plains, savannas, ice plains, and deserts), agent actions, and interactions. For real-world scenes, we utilize RealEstate10K (zhou2018stereo) with camera pose annotations to evaluate long-term world consistency.
Metrics. For quantitative evaluation, we employ reconstruction metrics, where the method of obtaining ground truth (GT) varies by specific settings. We then assess the consistency and quality of the generated videos using PSNR, LPIPS (zhang2018unreasonable), and reconstruction FID (rFID) (heusel2017gans), which collectively measure pixel-level fidelity, perceptual similarity, and overall realism.
Experimental details. For our experiments on Minecraft (fan2022minedojo), we utilize the Oasis (oasis2024) as the base model. Our model is trained using the Adam optimizer with a fixed learning rate of $2\times 10^{-5}$. Training is conducted at a resolution of $640\times 360$, where frames are first encoded into a latent space via a VAE at a resolution of $32\times 18$, then further patchified to $16\times 9$. Our training dataset comprises approximately 12K long videos, each containing 1500 frames, generated from fan2022minedojo. During training, we employ an 8-frame temporal context window alongside an 8-frame memory window. The model is trained for approximately 500K steps using 4 GPUs, with a batch size of 4 per GPU. For the hyperparameters specified in Algorithm 1 of the main paper, we set the similarity threshold $tr$ to 0.9, $w_{o}$ to 1, and $w_{t}$ to $0.2/t_{c}$. For the noise levels in Eq. (5) and Eq. (6), we set $k_{\text{min}}$ to 15 and $k_{\text{max}}$ to 1000.
For our experiments on RealEstate10K (zhou2018stereo), we adopt DFoT (song2025history) as the base model. The RealEstate10K dataset provides a training set of approximately 65K short video clips. Training is conducted at a resolution of $256\times 256$, with frames patchified to $128\times 128$. The model is trained for approximately 50K steps using 4 GPUs, with a batch size of 8 per GPU.
Figure 6: Results on RealEstate (zhou2018stereo). We visualize loop closure consistency over a full camera rotation. The visual similarity between the first and last frames serves as a qualitative indicator of 3D spatial consistency.
Table 4: Evaluation on RealEstate10K
Methods PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
CameraCtrl (he2024cameractrl) 13.19 0.3328 133.81
TrajAttn (xiao2024trajectory) 14.22 0.3698 128.36
Viewcrafter (yu2024viewcrafter) 21.72 0.1729 58.43
DFoT (song2025history) 16.42 0.2933 110.34
Ours 23.34 0.1672 43.14
@@ 4.1 Results on Generation Benchmark
Comparisons on Minecraft Benchmark. We compare our approach with a standard full-sequence (Full Seq.) training method (he2024cameractrl; wang2024motionctrl) and Diffusion Forcing (DF) (chen2025diffusion). The key differences are as follows: the full-sequence conditional diffusion transformer (peebles2023scalable) maintains the same noise level during training and inference, DF introduces different noise levels for training and inference, and our method incorporates a memory mechanism. To assess both short-term and long-term world consistency, we conduct evaluations within and beyond the context window. We evaluate both settings on 300 test videos. In the following experiments, the agent’s poses are generated by the game simulator as ground truth. However, in real-world scenarios, only the action input is available, and the pose is not directly observable. In such cases, the next-frame pose can be predicted based on the previous scenes, past states, and the upcoming action. We explore this design choice in the supplementary material.
@@
Figure 3 showcases WorldMem’s capabilities. The top section demonstrates its ability to operate in a free action space across diverse environments. Given a 600-frame memory bank, our model generates 100 future frames while preserving the ground truth’s actions and poses, ensuring strong world consistency. The bottom section highlights dynamic environment interaction. By using timestamps as embeddings, the model remembers environmental changes and captures natural event evolution, such as plant growth over time.
Comparisons on Real Scenarios. We compare our method with prior works (he2024cameractrl; xiao2024trajectory; yu2024viewcrafter; song2025history) on the RealEstate10K dataset (zhou2018stereo). We design 5 evaluation trajectories, each starting and ending at the same pose, across 100 scenes. The trajectory lengths range from 37 to 60 frames – exceeding the training lengths of all baselines (maximum 25 frames).
Table 5: Ablation on sampling strategy for training
Sampling strategy PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Small-range 13.23 0.3786 46.55
Large-range 15.11 0.3855 42.96
Progressive 19.32 0.1429 15.37
CameraCtrl (he2024cameractrl), TrajAttn (xiao2024trajectory), and DFoT (song2025history) discard past frames and suffer from inconsistency. Viewcrafter (yu2024viewcrafter) incorporates explicit 3D reconstruction, yielding better results, but is constrained by errors in post-processing such as reconstruction and rendering. As shown in Table 4 and Figure 6, our approach achieves superior performance across all metrics. However, the RealEstate dataset inherently limits the full potential of our method, as it consists of short, non-interactive clips with limited temporal complexity. We leave evaluation under more challenging and interactive real-world scenarios for future work.
@@ 4.2 Ablation
Time condition PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
w/o 17.17 0.1989 23.89
w/ 19.12 0.1613 16.53
Embedding designs. The design of embeddings within the memory block is crucial for cross-frame relationship modeling. We evaluate three strategies (Table 2): (1) sparse pose embedding with absolute encoding, (2) dense pose embedding with absolute encoding, and (3) dense pose embedding with relative encoding. Results show that dense pose embeddings (Plücker embedding) significantly enhance all metrics, emphasizing the benefits of richer pose representations. Switching from absolute to relative encoding further improves performance, particularly in LPIPS and rFID, by facilitating relationship reasoning and information retrieval. As illustrated in Figure 4.2, absolute embeddings accumulate errors over time, while relative embeddings maintain stability even beyond 300 frames.
Sampling strategy for training. We compare different sampling strategies during training in the Minecraft benchmark. Small-range sampling restricts memory conditioning to frames within 2m in the Minecraft world, while large-range sampling extends this range to 8m. Progressive sampling, on the other hand, begins with small-range samples for initial training steps and then gradually expands to large-range samples.
As shown in Table 5, both small-range and large-range sampling struggle with consistency and quality, whereas progressive sampling significantly improves all metrics. This suggests that gradually increasing difficulty during training helps the model learn to reason and effectively query information from memory blocks.
Time condition. We ablate the effectiveness of the timestamp condition (for both embedding and retrieval) in Table 6. We curate 100 video samples featuring placing events and evaluate whether future generations align with event progression. As shown in the table, incorporating the time condition significantly improves PSNR and LPIPS, indicating that adding temporal information helps the model faithfully reproduce event changes in world simulation. Since events like plant growth are inherently unpredictable, we do not conduct quantitative evaluations on such cases but instead provide qualitative illustrations in Figure 8.
Memory retrieve strategy. We analyze memory retrieval strategies in Table 3. Random sampling from the memory bank leads to poor performance and severe quality degradation, as evidenced by a sharp drop in rFID and rapid divergence from the ground truth (Figure 4.2). The confidence-based filtering significantly enhances consistency and generation quality. Additionally, we refine retrieval by filtering out redundant memory units based on similarity, further improving all evaluation metrics and demonstrating the effectiveness of our approach.
5 Limitations and Future works
Despite the effectiveness of our approach, certain issues warrant further exploration. First, we cannot guarantee that we can always retrieve all necessary information from the memory bank In some corner cases (e.g. , when views are blocked by obstacles), relying solely on view overlap may be insufficient. Second, our current interaction with the environment lacks diversity and realism. In future work, we plan to extend our models to real-world scenarios with more realistic and varied interactions. Lastly, our memory design still entails linearly increasing memory usage, which may impose limitations when handling extremely long sequences.
@@ 6 Conclusion
In conclusion, WorldMem tackles the longstanding challenge of maintaining long-term consistency in world simulation by employing a memory bank of past frames and associated states. Its memory attention mechanism enables accurate reconstruction of previously observed scenes, even under large viewpoints or temporal gaps, and effectively models dynamic changes over time. Extensive experiments in both virtual and real settings confirm WorldMem’s capacity for robust, immersive world simulation. We hope our work will encourage further research on the design and applications of memory-based world simulators.
Acknowledgements. This research is supported by the National Research Foundation, Singapore, under its NRF Fellowship Award <NRF-NRFF16-2024-0003>. This research is also supported by NTU SUG-NAP, as well as cash and in-kind funding from NTU S-Lab and industry partner(s).
7 Supplementary Materials
7.1 Details and Experiments
Embedding designs. We present the detailed designs of embeddings for timesteps, actions, poses, and timestamps in Figure 10, where $F,C,H,W,A$ denote the frame number, channel count, height, width, and action count, respectively.
The input pose is parameterized by position $(x,z,y)$ and orientation (pitch $\theta$ and yaw $\phi$). The extrinsic matrix $\mathbf{T}\in\mathbb{R}^{4\times 4}$ is formed as:
$\mathbf{T}=\begin{bmatrix}\mathbf{R}_{c}&\mathbf{c}\\ \mathbf{0}^{T}&1\end{bmatrix}$, (7)
where $\mathbf{c}=(x,z,y)^{T}$ and $\mathbf{R}_{c}=\mathbf{R}_{y}(\phi)\mathbf{R}_{x}(\theta)$.
To encode camera pose, we adopt the Plücker embedding. Given a pixel $(u,v)$ with normalized camera coordinates:
$\boldsymbol{\pi}_{uv}=\mathbf{K}^{-1}[u,v,1]^{T},$ (8)
its world direction is:
$\mathbf{d}_{uv}=\mathbf{R}_{c}\boldsymbol{\pi}_{uv}+\mathbf{c}.$ (9)
The Plücker embedding is:
$\mathbf{l}_{uv}=(\mathbf{c}\times\mathbf{d}_{uv},\mathbf{d}_{uv})\in\mathbb{R}^{6}.$ (10)
For a frame of size $H\times W$, the full embedding is:
$\mathbf{L}_{i}\in\mathbb{R}^{H\times W\times 6}.$ (11)
Memory context length. We evaluate how different memory context lengths affect performance in the Minecraft benchmark. Table 7 shows that increasing the context length from 1 to 8 steadily boosts PSNR, lowers LPIPS, and reduces rFID. However, extending the length to 16 deteriorates results, indicating that excessive memory frames may introduce noise or reduce retrieval precision. A context length of 8 provides the best trade-off, yielding the highest PSNR and the lowest LPIPS and rFID.
Pose prediction. For interactive play, ground truth poses are not accessible. To address this, we designed a lightweight pose prediction module that estimates the pose of the next frame. As illustrated in Figure 9, the predictor takes the previous image, the previous pose, and the upcoming action as inputs and outputs the predicted next pose. This module enables the system to operate using actions alone, eliminating the need for ground truth poses during inference. In Table 8, we compare the performance of using predicted poses versus ground truth poses. While using ground truth poses yields better results across all metrics, the performance drop with predicted poses is acceptable. This is because our method does not rely heavily on precise pose predictions – new frames are generated based on these predictions – and the ground truth poses generated by the Minecraft simulator also contain a certain degree of randomness.
Table 7: Ablation on length of memory context length
Length PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
1 16.18 0.1899 20.47
4 18.68 0.1568 16.54
8 19.32 0.1429 15.37
16 17.14 0.1687 18.33
Table 8: Comparison between using predicted poses and ground truth poses
Pose Type PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Ground truth 19.32 0.1429 15.37
Predicted 17.13 0.1786 20.36
Figure 9: Structure of pose predictor.
Figure 10: Illustration of different embeddings.
7.2 Memory Usage and Scalability Analysis
To assess the scalability and practical feasibility of our method, we provide detailed quantitative analysis covering memory usage, generation duration, training cost, and inference efficiency.
Memory Usage of the Memory Bank.
The memory bank is lightweight. Storing 600 visual memory tokens with shape $[600,16,18,32]$ in float32 takes approximately 21MB.
Retrieval Latency.
Below we report the average retrieval time (for 8 memory frames) as a function of memory bank size:
Number of Memory Candidates Retrieval Time (s)
10 0.04
100 0.06
600 0.10
1000 0.16
The generation cost (20 denoising steps) is $\sim$0.9s per frame. Retrieval time accounts for only 10–20% of total inference time even with 1000 candidates.
Comparison with Baseline.
We compare our method with a baseline model (without memory), under consistent settings: 8 context frames, 8 memory frames, 20 denoising steps, and no acceleration techniques, on single H200.
Training Inference
Method Mem. Usage Speed (it/s) Mem. Usage Speed (it/s)
w/o Memory 33 GB 3.19 9 GB 1.03
with Memory 51 GB 1.76 11 GB 0.89
Adding memory introduces moderate training overhead. During inference, the impact is minimal: only a small increase in memory usage and a slight decrease in speed.
Inference Optimization.
With modern acceleration techniques (e.g., timestep distillation, early exit, sparse attention), inference speed can reach $\sim$10 FPS, making our method practical for deployment.
FOV Overlapping Computation. We present the details of Monte Carlo-based FOV overlapping computation in Alg. 11, as well as the two-view overlapping sampling in Figure 11.
Input:
  • $Q_{\mathrm{ref}}\in\mathbb{R}^{F\times 5}$: reference poses from memory bank (x,y,z,pitch,yaw), $F$ is the number of stored poses.
  • $Q_{\mathrm{tgt}}\in\mathbb{R}^{5}$: pose of the current (target) frame.
  • $M$: number of 3D sample points (default 10,000).
  • $R$: radius of the sampling sphere (default 30 m).
  • $\phi_{h},\,\phi_{v}$: horizontal/vertical field-of-view angles (in degrees).
Output:
  • $\boldsymbol{\rho}\in\mathbb{R}^{F}$: overlapping ratios between each reference pose and the target pose.
begin
 $\Delta$ Step 1: Random Sampling in a Sphere
   Generate $M$ points $\mathbf{q}$ uniformly in a 3D sphere of radius $R$:
$\mathbf{q}\,\leftarrow\,\mathrm{PointSampling}\,\bigl(M,\,R\bigr).$
 $\Delta$ Step 2: Translate Points to $Q_{\mathrm{tgt}}$ as Center
   Let $Q_{\mathrm{tgt}}(x,y,z)$ be the 3D coordinates of the current camera pose. Shift all sampled points:
$\mathbf{q}\,\leftarrow\,\mathbf{q}\,+\,Q_{\mathrm{tgt}}(x,y,z).$
 $\Delta$ Step 3: FOV Checks
   Compute a boolean matrix $\mathbf{v}_{\mathrm{ref}}\in\{0,1\}^{F\times M}$, where each entry indicates if a point in $\mathbf{q}$ lies in the FOV of a reference pose:
$\mathbf{v}_{\mathrm{ref}}\,\leftarrow\,\mathrm{IsInsideFOV}\!\Bigl(\mathbf{q},\,Q_{\mathrm{ref}},\,\phi_{h},\,\phi_{v}\Bigr).$
Similarly, compute a boolean vector $\mathbf{v}_{\mathrm{tgt}}\in\{0,1\}^{M}$ for the target pose:
$\mathbf{v}_{\mathrm{tgt}}\,\leftarrow\,\mathrm{IsInsideFOV}\!\Bigl(\mathbf{q},\,Q_{\mathrm{tgt}},\,\phi_{h},\,\phi_{v}\Bigr).$
 $\Delta$ Step 4: Overlapping Ratio Computation
   Obtain the final overlapping ratio vector $\boldsymbol{\rho}\in\mathbb{R}^{F}$ by combining $\mathbf{v}_{\mathrm{ref}}$ and $\mathbf{v}_{\mathrm{tgt}}$. For instance,
$\boldsymbol{\rho}[i]\,=\,\frac{1}{M}\sum_{j=1}^{M}\Bigl(\mathbf{v}_{\mathrm{ref}}[i,j]\cdot\mathbf{v}_{\mathrm{tgt}}[j]\Bigr),$
to measure the fraction of sampled points that are visible in both the $i$-th reference pose and the target pose.
 Return $\boldsymbol{\rho}$.
end
Algorithm 2 Monte Carlo-based FOV Overlap Computation (Notationally Disjoint)
Figure 11: Two-view FOV overlapping visualization.
7.3 Visualizations
In this section, we provide more visualization of different aspects to facilitate understanding.
Minecraft Training Examples. We present a diverse set of training environments that include various terrain types, action spaces, and weather conditions, as shown in Figure 12. These variations help enhance the model’s adaptability and robustness in different scenarios.
Figure 12: Training Examples. Our training environments encompass diverse terrains, action spaces, and weather conditions, providing a comprehensive setting for learning.
Trajectory Examples in Minecraft. Figure 13 illustrates trajectory examples in the x-z space over 100 frames. The agent’s movement exhibits a random action pattern, ensuring diverse learning objectives and a broad range of sampled experiences.
Figure 13: Visualization of Trajectory Examples in the X-Z Space. The axis scales represent distances within the Minecraft environment.
Pose Distribution. We collect and visualize 800 samples within a sampling range of 8, as shown in Figure 14. The random pattern observed in Figure 14 ensures a diverse distribution of sampled poses in space, which is beneficial for learning the reasoning process within the memory blocks.
Figure 14: Visualization of Relative Pose Distribution for Training in X-Z Space. Red dots indicate positions, while yellow arrows represent directions.
More Qualitative Results. For additional qualitative examples, we recommend consulting the attached web page, which offers enhanced visualizations.